Circular No. 8954

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 L3 (HILL) R. E. Hill reports his discovery of another comet on CCD images taken

with the 0.68-m Catalina Schmidt telescope (discovery observation tabulated below), the co-added images from the discovery night showing a diffuse coma of diameter ∼ 10′′–12′′ and a diffuse, fan-shaped tail 1′–2′ long in p.a. 100^{o}; additional co-added images from June 14.2 UT show a poorly defined diffuse nuclear region, a coma diameter of 8′′–10′′, and a broad tail fan ∼ 2′ long in p.a. 110^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, several other CCD astrometrists have reported cometary appearance. J. E. McGaha (Tucson, AZ, U.S.A., 0.36-m reflector) writes that his stacked images from June 13.4 UT show a round 20′′ coma without condensation and an apparent ill-defined 45′′ elongation in p.a. 150^{o}; his stacked images from June 14.2 show a round 30′′ coma without condensation. R. Ligustri (Talmassons, Udine, Italy, 0.35-m reflector, June 13.9) notes that stacked images show a 30′′ coma elongated in p.a. 119^{o}. M. Pietschnig (Vienna, Austria, 0.35-m reflector, June 13.9) finds the comet diffuse with a 10′′ coma, a bright knot 5′′ from the coma’s center in p.a. 117^{o}, and a tail ∼ 30′′ long in p.a. 120^{o}. G. Hug (Eskridge, KS, U.S.A., 0.7-m reflector, June 14.14–14.15) detects a short, diffuse tail in p.a. ∼ 100^{o}. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m reflector, June 15.0) reports a diffuse, poorly condensed coma of diameter 10′′ in poor seeing.

2008 UT  α_{2000}  δ_{2000}  Mag.
June 13.27522  13^{h}50^{m}13.^{s}13  +58^{o}24′51′′.1  18.9

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-L56.

T = 2008 Apr. 20.148 TT  ω = 100.^{o}558
Ω = 23.885 }2000.0
q = 2.00238 AU  i = 100.404

COMET C/2008 A1 (McNAUGHT) Visual total-magnitude and coma-diameter estimates: Apr. 11.96 UT,

12.2, 2′ (M. Goiato, Araa̧tuba, Brazil, 22-cm reflector; moonlight); 24.94, 11.6, 4′ (Goiato); May 5.40, 11.8, 2′ (D. A. J. Seargent, Cowra, NSW, Australia, 25-cm reflector); 7.44, 11.4, 2′ (Seargent); 11.93, 10.8, 4′ (Goiato; moonlight); 25.92, 10.3, 5′ (Goiato); June 6.92, 9.9, 5′ (Goiato).

2008 June 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green